Wallace and Simpson formed a partnership with Wallace contributing $62,000 and Simpson contributing $42,000. Their partnership agreement calls for the income (loss) division to be based on the ratio of capital investments. The partnership had income of $130,000 for its first year of operation. When the Income Summary is closed, the journal entry to allocate partner income is: (Do not round intermediate calculations.)
Answer #1
Total capital investment 104000 =62000+42000
Net income allocated to Wallace 77500 =130000*62000/104000
Net income allocated to Simpson 52500 =130000*42000/104000
When the Income Summary is closed, the journal entry to allocate partner income is
Debit Income Summary $130,000; credit Wallace, Capital $77,500; credit Simpson, Capital $52,500.